# Eigenvalue Tracking of Large-Scale Systems Impacted by Time Delays

### v1

@@ Abstract
The paper focuses on tracking eigenvalue trajectories in power system models with time delays. We formulate a continuation-based approach that employs numerical integration to follow eigenvalues as system parameters vary, in the presence of one or multiple delayed variables. The formulation preserves the sparsity of the delay differential-algebraic equation (DDAE) system model and allows treating the delay magnitude itself as a varying parameter with implementation aspects discussed in detail. Accuracy is demonstrated on a modified IEEE 39-bus system with distributed energy resources. Scalability is discussed using a realistic dynamic model of the Irish transmission network.
[…]
This work is supported by Research Ireland (SFI), under grant number 21/SPP/3756 through the NexSys Strategic Partnership Programme.
@@ I-A Motivation
Time delays, arising from the growing volume of data processing and communication in inverter-dominated grids, are an increasingly important factor influencing power system dynamics. In particular, when present in closed-loop control settings, time delays can weaken damping and reduce the system’s stability margin [1, 2, 3, 4, 5]. These effects can be readily assessed using small-signal stability analysis (SSSA), where eigenvalue analysis plays a central role [6]. However, when delays are present, eigenvalue computations and tracking become significantly more complex, particularly in large-scale systems. This paper focuses on continuation-based methods for tracking eigenvalues in power system models with time delays.
@@ I-B Literature Review
Power system dynamics are commonly modeled through a set of differential-algebraic equations [7]. Introducing time delays transforms the model into delay differential-algebraic equations (DDAEs). In the context of SSSA, the stability of linearized DAEs and DDAEs is determined by their eigenvalues, defined as the roots of the system’s characteristic equation. For DDAEs, this equation is transcendental and gives rise to infinitely many eigenvalues [6]. Approximation techniques such as Padé polynomials or spectral discretization yield a finite-dimensional linear eigenvalue problem, which can then be solved using standard numerical algorithms [8]. These include dense-matrix methods (e.g., QR or QZ) for small to medium-sized systems, and sparse techniques (e.g., Krylov) for large-scale cases. However, when tracking eigenvalue trajectories as system parameters vary, repeated use of these solvers can be either computationally expensive or difficult to apply in a robust, automated manner.
@@
Continuation-based methods have been proposed as an alternative framework for tracking eigenvalue trajectories more efficiently. The core idea of continuation methods is to incrementally update the solution of a parameter-dependent problem using previously computed values. In the context of eigenvalue tracking, this solution consists of a subset of the system’s eigenpairs, i.e., eigenvalues with their associated eigenvectors. Setting aside time delays, [9] proposes a continuation-based algorithm that employs Newton’s method to follow eigenvalue trajectories of dynamical systems, highlighting the benefits of including a predictor step before each iteration. The authors in [10] combine this approach with a Cayley transform that updates the dimension of the considered subspace at every iteration, aiming to focus the computation on the modes most critical for stability. To evaluate stability margins of power systems, [11] derives eigenvalue and eigenvector sensitivities with respect to varying parameters and traces their trajectories through numerical integration. This approach makes it possible to identify poorly damped modes, in particular those that are critical for system stability.
@@
Despite the aforementioned advances in continuation-based tracking, only a limited number of works take into account time delays. Among them, [12] uses the derivative of the system’s characteristic equation with respect to system parameters to formulate a matrix equation of Sylvester type. The resulting tangent approximation allows tracking of eigenpairs, but the approach requires elimination of algebraic variables, restricting applicability to small and medium-sized power systems. In contrast, [13] retains the sparsity of the DDAE model. More precisely, [13] uses sensitivities with respect to delays. However, the incremental updates rely solely on Newton iterations, thus being sensitive to discontinuities and poor initialization (e.g., see [14]). Moreover, all existing works exclusively consider constant delays, effectively neglecting real-world communication network effects such as noise and data packet dropouts.
@@ I-C Contribution
This paper develops a continuation-based tracking formulation for power system models described by DDAEs. The formulation is general enough to permit multiple parameters to change simultaneously, as well as treating the delay magnitude as a varying parameter, enabling accurate estimation of stability margins. Extension to account for real-world communication network effects, such as noise and data packet dropouts, is duly discussed. These contributions are demonstrated through a comprehensive case study. First, the method is applied to systems with a single and multiple delays, tracking eigenvalues with respect to parameter variations. Second, the formulation considers the magnitude of the time delay as the varying parameter and its accuracy in capturing the stability margin is demonstrated. Finally, scalability and computational efficiency are assessed on a large-scale dynamic model of the all-island Irish transmission network.
@@ I-D Paper Organization
The remainder of the paper is organized as follows. Section II introduces preliminaries of SSSA. Section III discusses the formulation and implementation aspects of the proposed eigenvalue tracking approach. Section IV presents simulation results, based on the modified 39-bus system and a real-world scale model of the Irish transmission system, highlighting the validity of the method. Finally, Section V draws conclusions and outlines directions for future work.
@@ II-A DDAE Power System Model
Short-term power system stability in the presence of delays can be studied through a set of nonlinear DDAEs [6]:
@@
$\boldsymbol{x}^{j}_{d}=\boldsymbol{x}(t-\tau_{j})\ ,\quad\quad j=1,2,\ldots,\mu$ (3)
@@ II-B Eigenvalue Analysis
To study how the system responds to parameter variations, we introduce a scalar continuation parameter $p\in\mathbb{R}$. Then, (1) is rewritten as follows:
@@
where $\mbox{$\boldsymbol{\rm x}$}=(\Delta\boldsymbol{x},\Delta\boldsymbol{y})$, $\mbox{$\boldsymbol{\rm x}$}^{j}_{d}=(\Delta\boldsymbol{x}^{j}_{d},\Delta\boldsymbol{y}^{j}_{d})$ and
@@
with $\boldsymbol{\phi}$ being the associated right eigenvector.
@@ III Eigenvalue tracking
This section presents a continuation-based approach for tracking the eigenvalues of power systems with time delays. The method is first introduced for the case of a single delayed variable, then generalized to systems with multiple delays, and finally extended to allow the delay magnitude itself to vary as the continuation parameter.
@@ III-A Tracking in the Presence of a Single Delay
with $\dot{s}=\partial s/\partial p$, $\dot{\boldsymbol{\rm E}}=\partial{\boldsymbol{\rm E}}/\partial p$, $\dot{\boldsymbol{\rm A}}_{0}=\partial{\boldsymbol{\rm A}}_{0}/\partial p$, $\dot{\boldsymbol{\rm A}}_{1}=\partial{\boldsymbol{\rm A}}_{1}/\partial p$, $\dot{\boldsymbol{\rm\phi}}=\partial{\boldsymbol{\rm\phi}}/\partial p$. Equivalently, equation (12) is:
$(\boldsymbol{\rm E}+\tau\boldsymbol{\rm A}_{1}e^{-s\tau})\boldsymbol{\rm\phi}\dot{s}+(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{1}e^{-s\tau})\dot{\boldsymbol{\rm\phi}}=$ (14)
@@
Splitting real and imaginary parts, i.e., $\boldsymbol{\rm\phi}=\boldsymbol{\rm\phi}_{\rm r}+\jmath\boldsymbol{\rm\phi}_{\rm i}$, $s=s_{\rm r}+\jmath s_{\rm i}$, and by setting $\boldsymbol{\rm A}_{1}e^{-s_{\rm r}\tau}\cos{(s_{\rm i}\tau)}=\boldsymbol{\rm C}$, $\boldsymbol{\rm A}_{1}e^{-s_{\rm r}\tau}\sin{(s_{\rm i}\tau)}=\boldsymbol{\rm S}$, $\dot{\boldsymbol{\rm A}}_{1}e^{-s_{\rm r}\tau}\cos{(s_{\rm i}\tau)}=\boldsymbol{\rm C}_{\scriptscriptstyle D}$ and $\dot{\boldsymbol{\rm A}}_{1}e^{-s_{\rm r}\tau}\sin{(s_{\rm i}\tau)}=\boldsymbol{\rm S}_{\scriptscriptstyle D}$ we arrive at the following system:
@@
where $\boldsymbol{\rm y}=\boldsymbol{\rm y}(p)=(\boldsymbol{\rm\phi}_{\rm r},\boldsymbol{\rm\phi}_{\rm i},s_{\rm r},s_{\rm i})$, with $\boldsymbol{\rm y}\in\mathbb{R}^{2r+2}$; with:
@@ III-B Generalization to Multiple Delays
When multiple delays are present in the system, the matrix function $\boldsymbol{\rm P}$ is given by (10). Differentiating it with respect to the continuation parameter $p$ gives:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\sum_{j=1}^{\mu}(\dot{\boldsymbol{\rm A}}_{j}e^{-{s}\tau_{j}}-\tau_{j}\dot{s}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}})$ (19)
@@ III-C Delay Magnitude as Varying Parameter
In this subsection, we track the evolution of eigenvalues as the delay magnitude varies. We first consider the case where the system contains a single delayed variable, whose time delay has a varying magnitude and will be denoted by $\tau_{\ell}$. Its corresponding delayed state matrix is denoted by $\boldsymbol{\rm A}_{\ell}$. Given $p=\tau_{\ell}$, the matrix function $\boldsymbol{\rm P}$} is:
@@
Differentiation of (21) with respect to $p$ gives:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm A}}_{\ell}e^{-{s}p}+\boldsymbol{\rm A}_{\ell}(\dot{s}p+s)e^{-{s}p}$
In the above expression, $\boldsymbol{\rm A}_{\ell}$ does not explicitly depend on $\tau$, so $\dot{\boldsymbol{\rm A}}_{\ell}=0$. Substituting into (12) yields:
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-sp})\dot{\boldsymbol{\rm\phi}}+(\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}pe^{-sp})\boldsymbol{\rm\phi}\dot{s}$ $=$
$-s\dot{\boldsymbol{\rm E}}\,\boldsymbol{\rm\phi}+\dot{\boldsymbol{\rm A}}_{0}\boldsymbol{\rm\phi}-\boldsymbol{\rm A}_{\ell}e^{-sp}s$ $\boldsymbol{\rm\phi}$ (22)
Splitting real and imaginary parts of (22), we arrive at a system in the form of (18), where $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ are in this case functions of the varying delay magnitude $p=\tau_{\ell}$. Their definitions are detailed in the Appendix.
@@
We note that the term corresponding to $\tau_{\ell}$ remains separate, even if $\tau_{\ell}$ equals some $\tau_{j},\;j=1,2,\ldots,\mu$. Given that the delayed state matrices $\boldsymbol{\rm A}_{\ell}$ and $\boldsymbol{\rm A}_{j}$ do not explicitly depend on $\tau_{\ell}$, the derivative of (23) with respect to $p$ is:
$\dot{\boldsymbol{\rm P}}=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}+\sum_{j=1}^{\mu}\tau_{j}\dot{s}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}}+\boldsymbol{\rm A}_{\ell}(\dot{s}p+s)e^{-{s}p}$ (24)
@@
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-{s}p}-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}})\dot{\boldsymbol{\rm\phi}}$
$+(\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}\,p\,e^{-sp}+\sum_{j=1}^{\mu}\tau_{j}\boldsymbol{\rm A}_{j}e^{-s\tau_{j}})\boldsymbol{\rm\phi}\dot{s}=$
$-({s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm A}_{\ell}e^{-sp}s)\boldsymbol{\rm\phi}$ (25)
By splitting eigenvalue and eigenvector into real and imaginary parts we arrive at a system in the form of (18). The derivation of $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ is provided in the Appendix.

### v2

@@ Abstract
The paper focuses on tracking eigenvalue trajectories in power system models with time delays. We formulate a continuation-based approach that employs numerical integration to follow eigenvalues as system parameters vary, in the presence of one or multiple delayed variables. The formulation is compatible with sparse delay differential-algebraic equation (DDAE) formulations of the system model and allows treating the delay magnitude itself as a varying parameter with implementation aspects discussed in detail. The proposed approach is illustrated on a modified IEEE 39-bus system, as well as on a real-world-scale dynamic model of the Irish transmission network.
@@ I-A Motivation
Time delays, arising from the growing volume of data processing and communication in inverter-dominated grids, are an increasingly important factor influencing power system dynamics. In particular, when present in closed-loop control settings, time delays can weaken damping and reduce the system’s stability margin [9, 20, 2, 14, 17]. These effects can be readily assessed using small-signal stability analysis (SSSA), where eigenvalue analysis plays a central role [10]. However, when delays are present, eigenvalue computations and tracking become more involved, particularly in large-scale systems. This paper focuses on continuation-based methods for tracking eigenvalues in power system models with time delays.
@@ I-B Literature Review
Power system dynamics are commonly modeled through a set of differential-algebraic equations [5]. Introducing time delays transforms the model into delay differential-algebraic equations (DDAEs). In the context of SSSA, the stability of linearized DAEs and DDAEs is determined by their eigenvalues, defined as the roots of the system’s characteristic equation. For DDAEs, this equation is transcendental and gives rise to infinitely many eigenvalues [10]. Approximation techniques such as Padé polynomials or spectral discretization yield a finite-dimensional linear eigenvalue problem, which can then be solved using standard numerical algorithms [16]. These include dense-matrix methods (e.g., QR or QZ) for small to medium-sized systems, and sparse techniques (e.g., Krylov) for large-scale cases. However, when tracking eigenvalue trajectories as system parameters vary, repeated use of these solvers can be either computationally expensive or challenging to apply in a fully automated setting.
@@
Continuation-based methods have been proposed as an alternative framework for tracking eigenvalue trajectories more efficiently. The core idea of continuation methods is to incrementally update the solution of a parameter-dependent problem using previously computed values. In the context of eigenvalue tracking, this solution consists of a subset of the system’s eigenpairs, i.e., eigenvalues with their associated eigenvectors. For dynamical systems without time delays, [3] proposes a continuation-based algorithm that employs Newton’s method to follow eigenvalue trajectories, highlighting the benefits of including a predictor step before each iteration. The authors in [8] combine this approach with a Cayley transform that updates the dimension of the considered subspace at every iteration, focusing the computation on the modes most critical for stability. To evaluate stability margins of power systems, [18] derives eigenvalue and eigenvector sensitivities with respect to varying parameters and traces their trajectories through numerical integration. This approach makes it possible to identify poorly damped modes, in particular those that are critical for system stability.
@@
Despite the aforementioned advances in continuation-based tracking, only a very limited number of works have taken into account time delays. Among them, [12] uses the derivatives of the characteristic equation with respect to system parameters to formulate a Sylvester-type matrix equation. The resulting tangent approximation enables eigenpair tracking but requires elimination of algebraic variables, restricting applicability to small and medium-sized power systems. In contrast, [7] formulates eigen-sensitivities and tracking directly on the DDAE model, thereby retaining the sparsity structure. In that work, the incremental updates of the eigenpairs are obtained through Newton iterations, which may be sensitive to discontinuities and poor initialization (e.g., see [1]). Finally, existing works exclusively consider constant delays, effectively neglecting real-world communication network effects such as noise and data packet dropouts.
@@ I-C Contribution
This paper develops a continuation-based tracking formulation for power system models described by DDAEs. The formulation permits multiple parameters to vary simultaneously and allows the delay magnitude to be treated as a continuation parameter, enabling an accurate estimation of stability margins. Extension to account for real-world communication network effects, such as noise and data packet dropouts, is also discussed. These contributions are demonstrated through a comprehensive case study. First, the method is applied to systems with a single and multiple delays, tracking eigenvalues with respect to parameter variations. Second, the formulation considers the magnitude of the time delay as the varying parameter and its accuracy in capturing the stability margin is demonstrated. Finally, computational aspects are assessed on a large-scale dynamic model of the all-island Irish transmission network.
@@ I-D Paper Organization
The remainder of the paper is organized as follows. Section II introduces preliminaries on SSSA. Section III discusses the formulation and implementation aspects of the proposed eigenvalue tracking approach. Section IV presents simulation results, based on the modified 39-bus system and a real-world scale model of the Irish transmission system, highlighting the validity of the method. Finally, Section V draws conclusions and outlines directions for future work.
@@ II-A DDAE Power System Model
Short-term power system stability in the presence of delays can be studied through a set of nonlinear DDAEs in semi-implicit [^1] The semi-implicit form is often adopted because it can offer certain practical advantages, such as increased sparsity as well as flexibility in handling state and algebraic variables, see [10]. form [10]:
@@
$\boldsymbol{x}^{j}_{d}$ $=\boldsymbol{x}(t-\tau_{j})\ ,$ (3)
$\boldsymbol{y}^{j}_{d}$ $=\boldsymbol{y}(t-\tau_{j})\ ,$
where $j=1,2,\ldots,\mu$.
@@ II-B Eigenvalue Analysis
To study how the system responds to parameter variations, we introduce a scalar continuation parameter $p\in\mathbb{R}$, used to parametrize selected system quantities of interest. Then, (1) is rewritten as follows:
@@
where
$\boldsymbol{\rm x}$ $=(\Delta\boldsymbol{x},\Delta\boldsymbol{y})\ ,$
$\mbox{$\boldsymbol{\rm x}$}^{j}_{d}$ $=(\Delta\boldsymbol{x}^{j}_{d},\Delta\boldsymbol{y}^{j}_{d})\ ,$
and
@@
with $\boldsymbol{\phi}$ being the associated right eigenvector. The matrix function ${\boldsymbol{\rm P}}(s,p)$ plays an important role in the stability and dynamic behavior of (6).
@@ III Eigenvalue tracking
This section presents a continuation-based approach for tracking the eigenvalues of power systems with time delays. The method is first introduced for the case of a single delayed variable, then generalized to systems with multiple delays, and extended to allow the delay magnitude itself to vary as the continuation parameter. Handling stochastic, time-varying communication delays is also duly discussed.
@@ III-A Tracking in the Presence of a Single Delay
with:
$\dot{s}=\partial s/\partial p\ ,\quad\dot{\boldsymbol{\rm E}}$ $=\partial{\boldsymbol{\rm E}}/\partial p\ ,\quad\dot{\boldsymbol{\rm A}}_{0}=\partial{\boldsymbol{\rm A}}_{0}/\partial p\ ,$
$\dot{\boldsymbol{\rm A}}_{1}$ $=\partial{\boldsymbol{\rm A}}_{1}/\partial p\ ,\quad\dot{\boldsymbol{\rm\phi}}=\partial{\boldsymbol{\rm\phi}}/\partial p.$
Equivalently, equation (12) is:
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{1}e^{-s\tau})\dot{\boldsymbol{\rm\phi}}+(\boldsymbol{\rm E}+\tau\boldsymbol{\rm A}_{1}e^{-s\tau})\boldsymbol{\rm\phi}\dot{s}=$ (14)
@@
Splitting real and imaginary parts, i.e.
$\boldsymbol{\rm\phi}$ $=\boldsymbol{\rm\phi}_{\rm r}+\jmath\boldsymbol{\rm\phi}_{\rm i}\ ,$
$s$ $=s_{\rm r}+\jmath s_{\rm i}\ ,$
and by setting $\boldsymbol{\rm A}_{1}e^{-s_{\rm r}\tau}\cos{(s_{\rm i}\tau)}=\boldsymbol{\rm C}$, $\boldsymbol{\rm A}_{1}e^{-s_{\rm r}\tau}\sin{(s_{\rm i}\tau)}=\boldsymbol{\rm S}$, $\dot{\boldsymbol{\rm A}}_{1}e^{-s_{\rm r}\tau}\cos{(s_{\rm i}\tau)}=\boldsymbol{\rm C}_{\scriptscriptstyle D}$ and $\dot{\boldsymbol{\rm A}}_{1}e^{-s_{\rm r}\tau}\sin{(s_{\rm i}\tau)}=\boldsymbol{\rm S}_{\scriptscriptstyle D}$ we arrive at the following system:
@@
where $\boldsymbol{\rm y}=\boldsymbol{\rm y}(p)=(\boldsymbol{\rm\phi}_{\rm r},\boldsymbol{\rm\phi}_{\rm i},s_{\rm r},s_{\rm i})$, with $\boldsymbol{\rm y}\in\mathbb{R}^{2r+2}$; and:
@@ III-B Generalization to Multiple Delays
When multiple delays are present in the system, the expression (12) holds, where in this case the matrix function $\boldsymbol{\rm P}$ is given by (10). Differentiating (10) with respect to the continuation parameter $p$ gives:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\sum_{j=1}^{\mu}(\dot{\boldsymbol{\rm A}}_{j}-\tau_{j}\dot{s}\boldsymbol{\rm A}_{j})e^{-{s}\tau_{j}}$ (19)
@@ III-C Delay Magnitude as Varying Parameter
In this subsection, we track the evolution of eigenvalues as the delay magnitude is varied. We first consider the case where the system contains a single delayed variable, whose time delay has a varying magnitude and is denoted by $\tau_{\ell}$. The corresponding delayed state matrix is denoted by $\boldsymbol{\rm A}_{\ell}$. Given $p=\tau_{\ell}$, the matrix function $\boldsymbol{\rm P}$} is:
@@
Matrices $\boldsymbol{\rm E}$, $\boldsymbol{\rm A}_{0}$ and $\boldsymbol{\rm A}_{\ell}$ do not explicitly depend on $\tau$, hence differentiation of (21) with respect to $p$ yields:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}(\dot{s}p+s)e^{-{s}p}$
Substituting into (12) yields:
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-sp})\dot{\boldsymbol{\rm\phi}}+(\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}pe^{-sp})\boldsymbol{\rm\phi}\dot{s}=-\boldsymbol{\rm A}_{\ell}e^{-sp}s$ $\boldsymbol{\rm\phi}$ (22)
By splitting real and imaginary parts of (22), we arrive at a system in the form of (18), where $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ are in this case functions of the varying delay magnitude $p=\tau_{\ell}$. Their definitions are detailed in the Appendix.
@@
We note that the term corresponding to $\tau_{\ell}$ is kept separate, even if $\tau_{\ell}$ equals some $\tau_{j},\;j=1,2,\ldots,\mu$. Given that the matrices $\boldsymbol{\rm E}$, $\boldsymbol{\rm A}_{0}$, $\boldsymbol{\rm A}_{\ell}$ and $\boldsymbol{\rm A}_{j}$ do not explicitly depend on $\tau_{\ell}$, the derivative of (23) with respect to $p$ is:
$\dot{\boldsymbol{\rm P}}=\dot{s}\boldsymbol{\rm E}+\sum_{j=1}^{\mu}\tau_{j}\dot{s}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}}+\boldsymbol{\rm A}_{\ell}(\dot{s}p+s)e^{-{s}p}$ (24)
@@
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-{s}p}-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}})\dot{\boldsymbol{\rm\phi}}+$
$(\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}\,p\,e^{-sp}+\sum_{j=1}^{\mu}\tau_{j}\boldsymbol{\rm A}_{j}e^{-s\tau_{j}})\boldsymbol{\rm\phi}\dot{s}=-\boldsymbol{\rm A}_{\ell}e^{-sp}s\boldsymbol{\rm\phi}$ (25)
By splitting the eigenvalue and eigenvector into real and imaginary parts we arrive at a system in the form of (18). The derivation of $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ is provided in the Appendix.